I neeed help... I need the answer of some questions kindly help me out.

1 What is the function of neutron in the nucleus of the atom?
2 Is it possible to find the density of protons, neutrons and electrons or just the nuclei?
3 Electrons, as negatively charged particles, not only rotate about the core of atom but also about their own axis (spin) Why this happens?
4 A bird is perched on high voltage power line, why is the bird not electrocuted?
5 Two objects with different masses have the same kinetic energy. Which has the larger magnitude of momentum?
5Is it possible to find the density of protons, neutrons and electrons or just the nuclei?
1
Expert's answer
2019-04-24T10:10:25-0400

1 What is the function of neutron in the nucleus of the atom?

The neutrons and protons are made of quarks. Due to these quarks, neutrons and protons exchanges pi mesons which changes a proton into a neutron and vice versa. Without neutrons a nucleus can’t be stable.


2 Is it possible to find the density of protons, neutrons and electrons or just the nuclei?

The density of the nucleus is


"n=\frac {A} {\frac {4} {3} \pi R^3} (1)"

where A is the the mass number, R is the radius of nuclei

Using (1) we got: n=2.04×1017 kg×m3



3. Electrons, as negatively charged particles, not only rotate about the core of atom but also about their own axis (spin) Why this happens?

The “spin” of an electron or other fundamental particle is often described as “intrinsic angular momentum”. This terminology is required because a point particle or Planck length vibrating string cannot possess ½ ħ of angular momentum.


4 A bird is perched on high voltage power line, why is the bird not electrocuted?

The bird's body has a greater electrical resistance than the area between the limbs of the birds. Therefore, the electric current does not flow through the body of the birds.

 

5. Two objects with different masses have the same kinetic energy. Which has the larger magnitude of momentum?

The momentum of the heavier body is more.